In first case, two masses of 1 KG and 2KG are present on opposite sides{ends} of the frictionless pulley.


In second case, a mass of 1KG is on one side and 20 N force is applied on the other side of the pulley.


 What is the difference between the two situations?? PLS EXPLAIN where it matters.......


{TAKE g = 10 m/ sec2}


Grade:12

2 Answers

A S
13 years ago

In first case, net force along 2kg block is 20-10=10N. This force causes acceleration of total mass 1+2=3Kg. But, in second case, net force being same, total mass is 1 Kg only. Hence, acceleration is more is second case..... Here is the difference between two situations....

Aakash Anuj
13 years ago

See in the first case , the net force is 20-10  = 10 N and the net mass is 2+1 = 3 kg, Hence the acc. in the first case is 10/3 m/s^2.

In the second case the net force is again 20-10 - 10 N but here the net mass to be moved is only 1 kg . Hence the acc. in this case is 10/1 = 10 m/s^2